a man of weight 50 kg carries an object to height of 20m in 10 sec. the power used by man in this process is 4000w then find the weight of the object carried by man [assume g=10m/s
1
Expert's answer
2018-04-29T13:42:08-0400
power is calculated by the formula: P=A/t, where A - perfect job, J.; t – time, s.
In turn, the work will be calculated by the formula: A=F×h=m×g×h, where mass, m, is composed of their man weight and weight of cargo: m=m(man)+m(cargo), then:
P=((m(man)+m(cargo))×g×h)/t=4000 W.
P=((50+m(cargo))×10×20)/10=4000; 50+m(cargo)=200 kg; m(cargo)=200-50=150 kg.